OpenStudy (anonymous):

reduce the equation 3x + 6y -2 =0 to the slope-intercept form

OpenStudy (anonymous):

y=mx-b

OpenStudy (anonymous):

so 6y=-3x+2 y=-1/2x+1/3

OpenStudy (anonymous):

The formula is y=mx+b